# Dynamic Intelligence Assessment: Benchmarking LLMs on the Road to AGI with a Focus on Model Confidence

### v1

@@ Abstract
As machine intelligence evolves, the need to test and compare the problem-solving abilities of different AI models grows. However, current benchmarks are often overly simplistic, allowing models to perform uniformly well and, making it difficult to distinguish their capabilities. Additionally, benchmarks typically rely on static question-answer pairs, which models might memorize or guess. To address these limitations, we introduce the Dynamic Intelligence Assessment (DIA), a novel methodology for testing AI models using dynamic question templates and improved metrics across multiple disciplines such as mathematics, cryptography, cybersecurity, and computer science. The accompanying DIA-Bench dataset, which includes 150 diverse and challenging task templates with mutable parameters, is presented in various formats (text, PDFs, compiled binaries, and visual puzzles). Our framework introduces four new metrics to assess a model’s reliability and confidence across multiple attempts. These metrics revealed that even simple questions are frequently answered incorrectly when posed in varying forms, highlighting significant gaps in models’ reliability. Notably, models like GPT-4o tended to overestimate their mathematical abilities, while ChatGPT-4o demonstrated better decision-making and performance through effective tool usage. We evaluated eight state-of-the-art LLMs using DIA-Bench, showing that current models struggle with complex tasks and often display unexpectedly low confidence, even with simpler questions. The DIA framework sets a new standard for assessing not only problem-solving but also a model’s adaptive intelligence and ability to assess its own limitations. The dataset is publicly available on our project’s website. https://github.com/DIA-Bench
@@
component, formatting, style, styling, insert
@@ I Introduction
The introduction of neural networks and transformer-based architectures [4, 5, 6] enabled the development of powerful general-purpose models like BERT [7] and GPT-3 [8], which significantly advanced NLP capabilities. These models excelled across various tasks measured by benchmarks such as GLUE [9], SQuAD [10], and HumanEval [11]. However, these are static benchmarks, namely there is only one variant of each question and results could be memorized. Moreover, as noted by Wang et al. [12], many older benchmarks are no longer challenging enough to distinguish between newer models, as they often achieve near-perfect scores. For coding tasks, Honarvar et al. [13] explored the use of question templates to move beyond evaluating LLMs on separate, isolated problems. Apple recently extended the popular GSM8K dataset into a dynamic template-based benchmark [14]. While this allows for generating multiple variants of the same problems, the dataset remains limited to grade-school math and simple arithmetic tasks. Their findings show that LLMs struggle with even minor variations in these questions, relying more on pattern matching than true logical reasoning. In contrast, our work spans multiple disciplines and data formats, including cryptography, cybersecurity, and computer science. We move beyond accuracy-based metrics by introducing novel measures for assessing model reliability and confidence across repeated independent attempts, providing a more thorough and comprehensive evaluation of problem-solving consistency across diverse, real-world challenges.
Expanding on these, we propose Dynamic Intelligence Assessment (DIA), a testing framework that spans multiple disciplines, including cryptography, cybersecurity, and computer science. DIA moves beyond accuracy-based metrics by introducing novel measures for assessing model reliability and confidence across multiple independent instances generated from dynamic question templates. This approach allows the evaluation of LLMs consistency in problem-solving, under diverse, real-world challenges, offering a more comprehensive assessment without relying on in-context examples.
@@
  • RQ3: In problem-solving, does the use of tools impact the confidence and reliability of different models?
@@
    We introduce four key metrics for assessing performance on dynamic question templates. The Reliability Score measures a model’s performance on the dataset by penalizing incorrect answers. The Task Success Rate evaluates the consistency of correct answers for a given question template. The Confidence Index captures the percentage of question templates where all $k$ versions are correctly answered, and finally, the Near Miss Score counts the number of templates where $80\%\leq$ Task Success Rate$<100\%$ for a given $k$.
@@
    We introduce the Dynamic Intelligence Assessment (DIA) framework, to provide a comprehensive assessment of AI models’ reliability and confidence. This is done by novel evaluation metrics combined with DIA-Bench, a dataset of 150 diverse, hand-crafted dynamic question templates with varying difficulty.
@@
    We tested 8 popular state-of-the-art LLMs and ranked their problem-solving reliability and confidence, providing insights on which type of tasks are particularly challenging for current models.
In summary, our work redefines how AI models are assessed for reliability, by introducing both a dynamic testing methodology with four novel metrics, shifting the focus from one-off success to consistent, reliable and confident problem-solving. The paper is organized as follows: Section III details the methodology and dataset creation, Section IV details the experimental setup and results, Section II overviews related literature, Section V discusses limitations and ethical considerations, while Section VI concludes the paper.
@@ II-A Metrics for Benchmarking LLMs
Accuracy@k [8] measures the percentage of queries (from the entire test set) for which at least one relevant result was found within the top k results. Exact Match [10] demands an exact match with the expected output, making it particularly useful in tasks like question answering.
Robustness metrics evaluate a model’s performance under various challenging conditions [18], with Adversarial Robustness measuring the model’s resilience to attacks, such as handling unusual or malicious inputs [19]. Out-of-distribution robustness [19] measures the model’s performance on inputs significantly different from the training data, while False Refusal Rate (FRR) measures falsely rejecting benign prompts [20].
@@ II-B Benchmark Datasets
II-B1 Question Answering Benchmarks
@@
These are just a few examples, but in general, such tests suffer from the fact, that newer LLMs are getting better in lexical knowledge and thus these datasets while useful more measuring certain domain specific knowledge, no longer allow direct comparison between LLMs, if they all simply ace the test.
@@ II-B3 Security and Robustness Evaluation
In a small-scale study, Pearce et al. [42] demonstrated vulnerabilities in code generated by models like Copilot. The first large dataset to evaluate the security of AI-generated code was FormAI [43], Notably, it pioneered prompt templates with dynamically changing parameters to examine variations of similar programming questions. Turbulence [13] evaluates code generation with parameterized task templates but with a focus on task correctness. Wang et al. [12] introduced a framework to evolve existing datasets to be dynamic. Mirzadeh et al. [14] introduced a dynamic extension of GSM8K [44] to generate diverse variants of math problems. Their study indicated that LLMs struggle with variations in numerical values and irrelevant information, relying more on pattern matching than true logical reasoning, though the paper focuses solely on simple grade-school level arithmetic tasks. Our work expands dynamic testing by integrating diverse, dynamic, and increasingly complex challenges across multiple disciplines and data modalities.
@@ III Methodology
First, we introduce the four evaluation metrics, followed by a discussion of the dataset creation process.
Let $t$ denote a question template with mutable parameters, where $\mathcal{T}=\{t_{1},t_{2},\dots,t_{n}\}$ represents the set of such templates. For each $t\in\mathcal{T}$, the degree of freedom, denoted by $d(t)$, is the number of distinct questions that can be generated from a template. Furthermore, let $\mathcal{Q}(\mathcal{T},k)=\{q_{1},q_{2},\dots,q_{n\times k}\}$ denote the set of unique questions, where each template from $\mathcal{T}$ is used to generate $k$ different questions. In practice, we assume that $k\ll d(t)$. Let $\mathcal{S}_{\mathcal{Q}}=\{s_{1},s_{2},\dots,s_{n\times k}\}$ represent the set of solutions corresponding to $\mathcal{Q}$. To be more precise, there exists a mapping $f:\mathcal{Q}(\mathcal{T},k)\to\mathcal{S}_{\mathcal{Q}}$ such that $f(q_{i})=s_{i}$ for all $i\in\{1,2,\dots,n\times k\}$.
The Reliability Score measures the model’s performance across the dataset, with incorrect answers being heavily penalized, as shown in Formula 2. This approach is particularly useful in critical applications, where awareness of the tendency of hallucinations or incorrect responses is essential. By normalizing the score only by $k$, rather than by $k\times|\mathcal{T}|$, we can compare different instances of $\mathcal{Q}$, even when varying numbers of question instances are generated per template. Additionally, this approach still allows wrong answers to accumulate a significant negative score, providing a clear measure for human assessment about the model’s reliability over $\mathcal{T}$.
@@
The Reliability Score over a dataset $\mathcal{Q}(\mathcal{T},k)$ is calculated as:
@@
For a given question template $t_{i}$ with $k$ generated versions, Task Success Rate measures the number of correctly solved instances, where $i\in\{1,2,\dots,n\}$.
@@
represents the percentage of question templates in a dataset where, for a given template $t_{i}$, all $k$ generated queries are successfully answered,
@@
This metric is particularly useful in critical applications, such as critical infrastructure or autonomous vehicles, where flawless task execution is essential. As $k$ increases, the metric provides stronger reassurance of the model’s reliability.
A key statistic to consider is when a model solves most queries from a question template but misses a few.
@@
: counts the number of question templates where the model answers at least 80% but less than 100% of the instances correctly out of $k$.
@@
When we create multiple variations of the same core problem with slight adjustments, we assess the model’s ability to consistently solve these variations. We refer to this as Local Fuzzing, where task parameters are systematically modified to evaluate the robustness of the AI model.
Definition 5 (Local Task Fuzzing)
refers to the systematic process of generating a set of $k$ instances for $\forall t\in\mathcal{T}$, by varying the mutable parameters such that $k<d(t)$.
@@
DIA-Bench presents multi-modal challenges involving both visual and textual data, requiring models to process diverse formats such as PDFs and encoded files. It covers tasks including CAPTCHA solving, web security, reverse engineering, cryptography, mathematics, and logic. The multi-modal nature also stems from the fact that many questions demand specific tools like Python interpreters or Linux command-line utilities. These capabilities are not supported in LLM API calls as of today. Still, models should recognize when such tools are necessary and assess whether they can attempt a solution.
@@
An evaluation prompt consists of a general instruction template (Figure 2), and a dynamically generated challenge description, which also always ends with “Do NOT include any explanation or additional text. ALWAYS return the answer in this XML format: ’<xml>key</xml>’”. Examples for a challenge description can be seen in Figure 34, and 5. The combination of adhering to instructions while navigating complex tasks makes DIA-Bench particularly challenging. Research consistently demonstrates that LLMs face increasing difficulties as both the number and complexity of instructions grow [14, 45, 46, 47, 48].
Figure 2: General Part of the Prompt Template. (always included)
In our dataset, the difficulty of the tasks ranges from easy to highly challenging. Although including simpler tasks, such as finding the next prime number, may seem counterintuitive, they are particularly relevant for API calls in evaluating model confidence, a topic we explore further in our discussion.
@@
In our experiments, we tested ten Large Language Models, four of which were accessed via API calls, allowing for easy automation. We also included ChatGPT-4o, which, in some ways, acts as an orchestrator. It can run Python code, access the internet, operate within a Linux environment, and autonomously decide on the best solution. This makes it the closest candidate to general artificial intelligence as of today.
The verification module evaluates answers using an exact match, but it is flexible enough to accept minor variations, such as 5 being considered equivalent to 5.0. The evaluation process for ChatGPT-4o was time-consuming, as all queries had to be performed manually, since automating this process results in a ban from OpenAI. The 150 challenges are divided into two main categories: 50 Mathematics challenges and 100 Cybersecurity challenges. For this experiment, $k=5$ instances were generated from each of the 150 templates, resulting in a total of 750 unique questions. Figure 1 shows a visual overview of our methodology.
@@
Figure 3: Question Template 86. A CTF-style task where ChatGPT-4o and GPT-4o API both answered ’I do not know’ five times.
@@
Figure 4: Question Template 84: Reverse engineer. ChatGPT-4o and GPT-4o API both answered ’I do not know’ five times.
Figure 5: Question Template 76: Generate a SHA3-256 hash from the given plaintext message with adversarial prompting. GPT-4o API: Incorrect five times. ChatGPT-4o: Correct three times, Incorrect twice.
IV Discussion
@@
TABLE I: Performance of Tested Large Language Models with $k=5$ Instances Generated for Every Task Template. (Sorted by Reliability Score).
Evaluation Metrics
Model Tool Usage Correct Skipped Wrong Reliability Score Near Miss Score Confidence Index
ChatGPT-4o Yes 418 91 241 -64 14 38.67%
GPT-4o No 112 62 576 -208 3 10.6%
Gemini-1.5-pro No 80 0 670 -252 4 6.00%
Llama 3.1 No 65 0 685 -261 1 8.6%
Gemini-1.5-flash No 60 0 690 -264 0 4.67%
qwen2.5 No 62 4 684 -261.2 0 8.2%
yi No 1 0 749 -299.4 0 0
Mistral 7B No 0 0 750 -300 0 0
@@
We evaluated eight models: GPT-4o, Gemini-1.5-pro, Gemini-1.5-flash, Llama 3.1, qwen2.5, Mistral 7B, and yi, along with one chat model equipped with tool-using capabilities, ChatGPT-4o. Notably, existing literature often overlooks whether the LLMs under examination possess tool-usage capabilities, and many benchmarks fail to specify whether the chat or API version of a model is employed.
In this section, we focus primarily on comparing the OpenAI models. First, because they are the top performers, and second, one can use tools while the other cannot, offering the most compelling case for comparison. While the initial assumption might be that ChatGPT-4o performs better solely because of its tool-using capabilities, a more nuanced analysis suggests that the gap between the two models is not purely a function of tool availability.

### v3

@@ Abstract
As machine intelligence evolves, the need to test and compare the problem-solving abilities of different AI models grows. However, current benchmarks are often simplistic, allowing models to perform uniformly well and making it difficult to distinguish their capabilities. Additionally, benchmarks typically rely on static question-answer pairs that the models might memorize or guess. To address these limitations, we introduce Dynamic Intelligence Assessment (DIA), a novel methodology for testing AI models using dynamic question templates and improved metrics across multiple disciplines such as mathematics, cryptography, cybersecurity, and computer science. The accompanying dataset, DIA-Bench, contains a diverse collection of challenge templates with mutable parameters presented in various formats, including text, PDFs, compiled binaries, visual puzzles, and CTF-style cybersecurity challenges. Our framework introduces four new metrics to assess a model’s reliability and confidence across multiple attempts. These metrics revealed that even simple questions are frequently answered incorrectly when posed in varying forms, highlighting significant gaps in models’ reliability. Notably, API models like GPT-4o often overestimated their mathematical capabilities, while ChatGPT-4o demonstrated better performance due to effective tool usage. In self-assessment OpenAI’s o1-mini proved to have the best judgement on what tasks it should attempt to solve. We evaluated 25 state-of-the-art LLMs using DIA-Bench, showing that current models struggle with complex tasks and often display unexpectedly low confidence, even with simpler questions. The DIA framework sets a new standard for assessing not only problem-solving, but also a model’s adaptive intelligence and ability to assess its limitations. The dataset is publicly available on the project’s page: https://github.com/DIA-Bench.
@@
Artificial Intelligence, Large Language Models, Dynamic Benchmarking, Performance Metrics, Reliability
@@ I Introduction
The introduction of neural networks and transformer-based architectures [4, 5, 6] enabled the development of powerful general-purpose models like BERT [7] and GPT-3 [8], which significantly advanced NLP capabilities. These models excelled across various tasks measured by benchmarks such as GLUE [9], SQuAD [10], and HumanEval [11]. However, these are static benchmarks; namely, there is only one variant of each question, and results could be memorized. Moreover, as noted by Wang et al.[12], many older benchmarks are no longer challenging enough to distinguish between newer models, as they often achieve near-perfect scores. For coding tasks, Honarvar et al.[13] explored using question templates to move beyond evaluating LLMs on separate, isolated problems. Apple recently extended the popular GSM8K dataset into a dynamic template-based benchmark [14]. While this allows for generating multiple variants of the same problems, the dataset remains limited to grade-school math and simple arithmetic tasks. Their findings show that LLMs struggle with even minor variations in these questions, relying more on pattern matching than logical reasoning.
To address these challenges, we introduce DIA-Bench, a benchmark dataset of 150 question templates covering multiple disciplines—like mathematics, cybersecurity, or CTF challenges—and various data formats and modalities. Each template in DIA-Bench can produce multiple distinct questions with different answers, where the difficulty ranges from easy to extremely challenging questions, that as of today none of the examined models were able to solve.
We propose the Dynamic Intelligence Assessment (DIA) framework, a novel benchmarking approach that moves beyond traditional accuracy-based metrics. The DIA framework utilizes the dynamic questions templates, and introduces innovative measures for evaluating model reliability and confidence across multiple independent attempts on a given task type. By evaluating models’ consistent problem-solving abilities across diverse, real-world challenges without relying on in-context examples, the DIA framework offers a comprehensive assessment of model performance. This provides better insights into models’ confidence and reliability across a broad spectrum of tasks, enhancing our understanding of their true capabilities.
@@
  • RQ2: How can a benchmark dataset better accurately measure LLMs’ confidence and reliability in problem-solving?
  • RQ3: How does the use of tools impact the confidence and reliability of various models in problem-solving?
@@
    We introduce four key metrics as part of the DIA framework. The Reliability Score measures a model’s performance on the entire dataset by penalizing incorrect answers, and allowing a model to not answer if it is not certain. Task Success Rate counts the number of correct answers for all $k$ versions of a given question template. The Confidence Index captures the percentage of question templates where all $k$ versions are correctly answered. Finally, the Near Miss Score counts the number of question templates where the model correctly answers at least $80\%$ but less than $100\%$ of question instances.
@@
    We introduce DIA-Bench, a dataset of 150 diverse, hand-crafted dynamic question templates with varying difficulty.
@@
    We tested 25 popular state-of-the-art LLMs and ranked their problem-solving reliability and confidence, providing insights on which types of tasks are particularly challenging for current models.
In summary, our work redefines how AI models are assessed for reliability by introducing a dynamic testing methodology with four novel metrics, shifting the focus from one-off success to consistent, reliable, and confident problem-solving. The paper is organized as follows: Section II reviews the related literature and existing benchmarks evaluating AI capabilities. Section III details the methodology and dataset creation, Section IV details the experimental setup and results. Section V discusses limitations and ethical considerations, while Section VI concludes the paper.
@@ II-A Metrics for Benchmarking LLMs
Accuracy@k [8] measures the percentage of queries (from the entire test set) for which at least one relevant result was found within the top k results. Exact Match [10] requires the output to precisely match the expected result, making it especially valuable for tasks such as question answering.
Robustness metrics assess a model’s performance in diverse challenging scenarios [18], where Adversarial Robustness evaluates the model’s resilience to attacks, such as handling unusual or malicious inputs [19]. Out-of-distribution robustness [19] measures the model’s performance on inputs significantly different from the training data. False Refusal Rate (FRR) measures falsely rejecting benign prompts [20].
@@ II-B Benchmark Datasets
II-B1 Question-Answering Benchmarks
@@
In general, such tests suffer from the fact that newer LLMs are getting better in lexical knowledge, therefore these datasets while useful for measuring certain domain-specific knowledge, no longer allow direct comparison between LLMs, as they often perform uniformly well [22].
@@ II-B3 Security and Robustness Evaluation
In a small-scale study, Pearce et al. [42] demonstrated vulnerabilities in code generated by models like Copilot. The first large dataset to evaluate the security of AI-generated code was FormAI [43]. Notably, it pioneered prompt templates with dynamically changing parameters to examine variations of similar programming questions. Turbulence [13] evaluates code generation with parameterized task templates but with a focus on task correctness. Wang et al. [12] introduced a framework to evolve existing datasets to be dynamic. Mirzadeh et al. [14] introduced a dynamic extension of GSM8K [44] to generate diverse variants of math problems. Their study indicated that LLMs struggle with variations in numerical values and irrelevant information, relying more on pattern matching than true logical reasoning, though the paper focuses solely on simple grade-school level arithmetic tasks. Our work expands dynamic testing by integrating diverse, dynamic, and increasingly complex challenges across multiple disciplines and data modalities.
@@ III Methodology
Figure 1 shows a visual overview of our methodology. We begin by introducing the four evaluation metrics, followed by a detailed discussion of the dataset creation process.
Let $t$ denote a question template with mutable parameters, where $\mathcal{T}=\{t_{1},t_{2},\dots,t_{n}\}$ represents the set of such templates. For each $t\in\mathcal{T}$, the degree of freedom, denoted by $d(t)$, is the number of distinct questions that can be generated from a template. Furthermore, let $\mathcal{Q}(\mathcal{T},k)=\{q_{1},q_{2},\dots,q_{n\times k}\}$ denote the set of unique questions, where each template from $\mathcal{T}$ is used to generate $k$ different questions.
In practice, we aim for all generated questions from a template $t_{i}$ to be unique. If $d(t)$ is not sufficiently large, randomly altering the mutable parameters in $t_{i}$ might result in duplicate questions, which we want to avoid. To address this, a more strategic approach is required when creating the questions from templates to ensure the final dataset is diverse. We refer to this approach as Local Task Fuzzing, where parameters are systematically modified instead of relying on randomization.
Definition 1 (Local Task Fuzzing)
Local Task Fuzzing refers to the systematic process of generating a set of $k$ instances for $\forall t\in\mathcal{T}$, by varying the mutable parameters such that $k<d(t)$.
III-A Evaluation metrics
When a dataset $\mathcal{Q}(\mathcal{T},k)$ is constructed with diverse questions, a model’s performance on this dataset can be evaluated using various metrics. We propose four novel metrics to evaluate a model’s confidence and reliability.
Let $\mathcal{S}_{\mathcal{Q}}=\{s_{1},s_{2},\dots,s_{n\times k}\}$ represent the set of solutions corresponding to $\mathcal{Q}$. To be more precise, there exists a mapping $f:\mathcal{Q}(\mathcal{T},k)\to\mathcal{S}_{\mathcal{Q}}$ such that $f(q_{i})=s_{i}$ for all $i\in\{1,2,\dots,n\times k\}$.
@@ III-A Evaluation metrics
The Reliability Score (RS@k) over a dataset $\mathcal{Q}(\mathcal{T},k)$ is calculated as:
@@
RS@k measures the model’s performance across the dataset, with incorrect answers being heavily penalized, as shown in Formula 2. This approach is particularly useful in critical applications, where awareness of the tendency of hallucinations or incorrect responses is essential. Normalizing the score by $k$ enables comparisons across different instances of $\mathcal{Q}$, even when the number of question instances per template varies. However, normalizing the final score by $k\times|\mathcal{T}|$ would not accurately represent the true strength of this metric, as it is essential for the RS score to reflect the number of question templates used in the evaluation. Additionally, this approach allows wrong answers to accumulate a significant negative score, providing a clear measure for human assessment about the model’s reliability over $\mathcal{T}$.
@@
The Task Success Rate (TSR@$(t_{i},k)$) evaluates the number of correct answers for a given question template $t_{i}$ out of the $k$ generated instances, where $i\in\{1,2,\dots,n\}$.
@@
The Confidence Index (Conf@k) represents the percentage of question templates in a dataset where, for a given template $t_{i}$, all $k$ generated queries are successfully answered,
@@
This metric is particularly useful in critical applications, such as critical infrastructure or autonomous vehicles, where flawless task execution is essential. As $k$ increases, the metric provides stronger reassurance of the model’s reliability. A key statistic to consider is when a model solves most queries from a question template but misses a few.
@@
The Near Miss Score (NMS@k) counts the number of question templates where the model answers at least 80% but less than 100% of the instances correctly out of $k$.
@@ III-B Dataset and Prompt Engineering
DIA-Bench presents multi-modal challenges involving both visual and textual data, requiring models to process diverse data formats, including PDFs and encoded files. The dataset covers a wide range of tasks, such as CAPTCHA solving, cryptography, reverse engineering, web security, mathematics, and logic. Many questions require the use of tools, such as Python interpreters or Linux command-line utilities. Although these capabilities are not currently supported in LLM API calls, models should recognize when such tools are needed to provide an appropriate response. The tasks in our dataset range in difficulty from easy to highly challenging. Although including simpler tasks, such as finding the next prime number, may seem counterintuitive, they are particularly relevant for evaluating model reliability especially for models without tools. The next-prime problem for large numbers is a typical case where reasoning alone would be insufficient, and a tool (like running Python code) is required. Since guessing the next prime is impractical, a model should recognize that this task cannot be solved without the appropriate tool and should skip it if tool usage is unavailable.
@@
In DIA-Bench, a single prompt includes a general instruction template and a dynamically generated challenge description, as shown in Figure 2. Examples for a challenge description can be seen in Figure 34, and 5. The combination of adhering to instructions while navigating complex tasks makes DIA-Bench particularly challenging. Research consistently demonstrates that LLMs face difficulties as the number or complexity of instructions grow [14, 45, 46, 47, 48].
Figure 2: General Part of the Prompt Template.
@@
We tested $25$ Large Language Models, $24$ of which were accessed via API calls, allowing for easy automation. We also included ChatGPT-4o, which, in some ways, acts as an orchestrator. It can run Python code, access the internet, operate within a Linux environment, and autonomously decide on the best solution. This makes it the closest candidate to artificial general intelligence as of today. The evaluation process for ChatGPT-4o was time-consuming, as all queries had to be performed manually. The $150$ templates are divided into two main categories: $50$ Mathematics and $100$ Cybersecurity & Computer Science challenges.
@@
Figure 3: Question Template 86. A CTF-style task where ChatGPT-4o and GPT-4o API both answered ’I do not know’ five times. o1-mini: Correct 4 times, Incorrect 1.
@@
Figure 4: Question Template 84: Reverse engineering. ChatGPT-4o, GPT-4o API, and o1-mini API all answered ’I do not know’ five times.
Figure 5: Question Template 76: GPT-4o API: Incorrect 5. ChatGPT-4o: Correct 3, Incorrect 2. o1-mini: Incorrect 4, skipped 1
(a) Mathematics Questions
(b) Computer Science and Cybersecurity Questions
Figure 6: Performance of the Top 8 Examined Models in Different Categories
IV Discussion & Results
@@
TABLE I: Performance of Tested LLMs with $k=5$ Instances Generated for Every Task Template. (Sorted by Confidence Index)
Results Evaluation Metrics
Model Company Size License Tool* Correct Skipped Wrong RS NMS CI
ChatGPT-4o OpenAI N/A Proprietary 418 91 241 -64 14 38.67%
o1-mini OpenAI N/A Proprietary X 269 280 201 -26.6 14 21.34%
GPT-4o OpenAI N/A Proprietary X 182 70 498 -162.80 29 17.33%
GPT-4o-mini OpenAI N/A Proprietary X 123 29 598 -214.60 20 11.33%
LLama-3.1-Nemotron NVIDIA 70B llama3.1 X 109 21 620 -226.20 3 9.33%
Codegemma Google 7B gemma X 81 0 669 -251.4 0 8%
Mistral NeMo Mistral AI 12B Apache-2.0 X 84 15 651 -243.6 1 7.33%
Gemini-1.5-pro Google N/A Proprietary X 80 0 670 -252.0 4 6.00%
Llama3.1 Meta 8B llama3.1 X 65 0 685 -261.0 0 4.67%
Gemini-1.5-flash Google N/A Proprietary X 60 0 690 -264.0 4 4.67%
Dolphin-2.8 Mistral Cognitive 7B Apache-2.0 X 63 22 665 -253.4 5 4%
Mistral-openorca Microsoft 7B Apache-2.0 X 66 0 684 -260.4 4 4%
WizardLM2 Microsoft 7B Apache-2.0 X 57 1 692 -265.4 4 4%
CodeQwen1.5 Qwen 7B tongyi-qianwen X 65 22 663 -252.2 7 3.33%
Qwen2.5 Qwen 7B Apache-2.0 X 62 9 684 -261.20 3 3.33%
Mixtral-8x7B Mistral AI 47B Apache-2.0 X 60 0 690 -264.0 5 3.33%
Gemma Google 7B gemma X 53 0 697 -268.2 2 3.33%
llava-v1.5 liuhaotian 7B llama2 X 48 0 702 -271.2 2 2.67%
Phi3 Microsoft 3B MIT X 26 2 722 -283.6 0 2%
Codellama Meta 7B llama2 X 19 1 730 -288.2 0 0%
Deepseek-coder Depseek AI 33B deepseek X 19 1 730 -288.2 0 0%
Qwen2.5 Qwen 3B Apache-2.0 X 16 9 725 -286.8 0 0%
Orca-mini Microsoft 3B llama X 5 0 745 -297.0 0 0%
Wizard-Vicuna-Uncensored Cognitive 7B llama X 0 4 746 -298.4 0 0%
Llama2-Uncensored Cognitive 7B llama2 X 0 1 749 -299.6 0 0%
*Tool usage means that ChatGPT-4o can execute codes, access the internet, and run bash commands within a sandboxed environment.
$\dagger$ The o1-mini model attempted to skip 52 additional tasks but failed to provide the correct XML format. See Section V for detailed comments.
We evaluated 25 models, with the complete list available in Table I. Using $k=5$, we generated five different questions per template, resulting in a total of 750 questions. Higher $k$ values can provide an even more accurate assessment.
Some questions templates produce a particularly large prompt (e.g., containing a Base64-encoded image), therefore certain questions can exceed the token limit of what some models can handle. For example a full prompt from question template $135$ results in $17,667$ tokens  [^3] https://platform.openai.com/tokenizer.
@@
Among the tested models, currently only ChatGPT-4o has the capability to use tools. It can write and run code, access the web, as well as it has support for multimodal data like visual images, voice, and textual data. Notably, existing literature often overlooks whether the LLMs under examination possess tool-usage capabilities, and many benchmarks fail to specify whether the chat or API version of a model is employed.
This section examines common mistakes made by models and highlights model-specific weaknesses. Particular attention is given to OpenAI’s models, as they provide the best ground for comparing a tool enabled model (ChatGPT-4o) and the equivalent API model (GPT-4o) that lacks this functionality. While it might initially appear that ChatGPT-4o’s performance advantage stems solely from its ability to use tools, a closer analysis reveals that the performance gap between OpenAI’s models is influenced by factors beyond tool availability alone.